A bullet of mass a and velocity b is fired into a large block of mass c.  The final velocity of the system is? 

Grade:9

3 Answers

ABHIJITHU S
5 years ago
Applying conservation of linear momentum,
 Initial momentum is ab.
Final momentum is (a+c)*final velocity.
There for final velocity=ab/(a+c)[momentum before collision is equal to momentum after collision]
Khimraj
5 years ago
As the external force is zero so we can apply conservation of momentum law
So
a*b = (a+c)*v
Then v = ab/(a+c) m/s
Hope it clears. If you like answer then please approve it.
Kiran
5 years ago
Mass of bullet=a (m1)
initial velocity of bullet=b (u1)
Mass of large block=c (m2)
Initial velocity of large block=0 (u2)
M1v2 + m2u2= v(m1+m2)
(a*b)+(c*0)=v(a+c)
ab=v(a+c)
V=ab
    (a+c)